OpenStudy (anonymous):

he perimeter of a face of a cube is 16 mm. What is its volume? 64 mm3 128 mm3 512 mm3 4096 mm3

OpenStudy (anonymous):

one side = 16 / 4 = 4 volume = 4^3 = 64